OpenStudy (anonymous):

Harry is building a house with a sloping roof. If the roof is 44 feet long and 10 feet high, what is the angle that the roof makes with the horizontal? a) 12.8 degrees b) 13.1 degrees c) 14.6 degrees d) 15.3 degrees e) 17.5 degrees please help !

OpenStudy (anonymous):



OpenStudy (anonymous):

use tan. tanx = opp/adj tanx = 10/44 x = arctan 10/44 use a calculator to find the answer, good luck

OpenStudy (anonymous):

i dont have a calculator with me at the moment..

OpenStudy (anonymous):

use an online one.

OpenStudy (anonymous):

i got it wrong.. i got 0.9 :/